Art realized that he had 2 more quarters than he had originally thought in his pocket. If all of the change in his pocket is quarters and it totals to $8.75, how many quarters did he originally think were in his pocket?
Answer by askmemath(368)

Art had $8.75 or 875 cents in his pocket
This amount was in quarters
So,number of quarters = 875/25 = 35
From this Art thought he had 2 more than he orginally thought so he thought he had 35-2 = 33 .